# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.3 Part $M\subset M_{0}$
$\left\|\mathbf{III}^{4}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|$
$\leq T^{\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|^{2}$
×$\times\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}T^{-(1+\tau)}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
$=O_{p}(T^{\tau}),$ (A.27)
while
@@
$\left\|\mathbf{III}^{2}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|$
$\leq T^{\tau}\left\|\boldsymbol{\Gamma}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
×$\times\underbrace{\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|}_{O_{p}(N^{-1/2})}$
$=O_{p}(T^{\tau}N^{-1/2}),$ (A.28)
which is negligible only if $\tau\in(0,0.5)$ under $TN^{-1}=O(1)$ (to relate this term to (6.3) and the subsequent similar terms), and the same condition applies to $\mathbf{III}^{3}_{d}$:
@@
$\left\|\mathbf{III}^{3}_{d}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}\right\|$
@@
$\leq T^{\tau}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}\right\|$
×$\times\underbrace{\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\eta}_{\boldsymbol{\Gamma},i}\right\|}_{O_{p}(N^{-1/2})}$
$=O_{p}(T^{\tau}N^{-1/2}),$ (A.29)
as welll. Next, we move on to $\mathbf{III}_{a}$, and for this we firstly analyze
@@
$\left\|\mathbf{III}_{a}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
@@
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
@@
$\leq T^{-1}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|$
$=O_{p}(T^{-1}),$ (A.31)
@@
which would not be negligible under correlated idiosyncratics (see Lemma 2). The remaining two terms, are slowly vanishing. In particular,
@@
$\left\|\mathbf{III}_{b}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq T^{(\tau-1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
@@
×$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
$=O_{p}(T^{(\tau-1)/2}),$ (A.32)
and
@@
$\left\|\mathbf{III}_{c}\right\|$ $=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
@@
$=\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$

### v3

@@ 5.3.1 Case of $M\subset M_{0}$
$\mathrm{tr}\left(\mathbf{III}_{d}\right)=\mathrm{tr}\left(\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right)$
@@
$=\mathrm{tr}\left(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\left[\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}\right]\right)$
@@
$=\mathrm{tr}\left(\left[\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}\right]T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right)$
$\geq\lambda_{\min}\left(\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}\right)$
@@
$\times\mathrm{tr}\left(T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right)$
$\to_{p}c>0$ (A.27)
as $(N,T)\to\infty$, for a positive constant $c$ by the results in Fang et al., (1994). This follows from the fact that second positive semi-definite matrix in the product is non-zero by Exercise 8.26 (c) in Abadir and Magnus, (2005). Particularly, for a positive definite $\mathbf{A}\in\mathbb{R}^{p\times p}$ and $\mathbf{B}\in\mathbb{R}^{p\times q}$, $\mathbf{B}^{\prime}\mathbf{A}\mathbf{B}=\mathbf{0}_{q\times q}$ if and only if $\mathbf{B}=\mathbf{0}_{p\times q}$, which means that $\mathrm{rk}(\mathbf{B})=0$. In our case, $\mathbf{B}=T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}$ and we know that $\mathrm{rk}\left(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right)=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}{\mathbf{q}}_{M^{C}})=m_{x}-g>0$ even in the limit. Also, $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}\boldsymbol{\Gamma}_{i}^{\prime}$ is an $m_{x}\times m_{x}$ positive definite matrix by our assumptions.
Next in our analysis, we move on to $\mathbf{III}_{a}$, and for this we firstly analyze
@@
$\left\|\mathbf{III}_{a}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
@@
$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$
@@
$\leq T^{-1-\tau}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|^{2}\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}\right\|^{2}\left\|(T^{-(1+\tau)}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|$
$=O_{p}(T^{-1-\tau})=o_{p}(T^{-\tau}),$ (A.29)
@@
which would not be negligible under correlated idiosyncratics (see Lemma 2). The remaining two terms are vanishing, as well. In particular,
@@
$\left\|\mathbf{III}_{b}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}\right\|$
@@
$\leq T^{-(\tau+1)/2}\frac{1}{N}\sum_{i=1}^{N}\left\|T^{-(1+\tau)/2}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|\left\|\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}\right\|^{2}\left\|T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\right\|$
@@
$\times\left\|(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}T^{-(1+\tau)}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\right\|\left\|\boldsymbol{\Gamma}_{i}\right\|$
$=O_{p}(T^{-(\tau+1)/2})=o_{p}(T^{-\tau}),$ (A.30)
for $\tau\in(0,1)$, and
@@
$\left\|\mathbf{III}_{c}\right\|$ $=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{V}_{i}\right\|$
@@
$=\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\Gamma}_{i}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}(\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}})^{+}\mathbf{q}_{M^{C}}^{\prime}\overline{\boldsymbol{\Gamma}}^{\prime}\mathbf{F}_{\mathbf{x}}^{\prime}\mathbf{M}_{\mathbf{F}_{\mathbf{x}}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{V}_{i}\right\|$